OpenStudy (abbydale74):

A megabyte is two to the twentieth bytes. A terabyte is two to the fortieth bytes. Using the exponent rule for division, find how many megabytes are in a terabyte?

OpenStudy (agreene):

\[\text{MB}=2^{20}\\ \text{TB}=2^{40}\\\frac{\text{MB}}{\text{TB}}=\frac{2^{20}}{2^{40}}\] is what this is asking, what are your thoughts on the solution?

OpenStudy (texaschic101):

look at this and see if you can figure it out...this is an exponent rule.. (x^m) / (x^n) = x^(m - n)